the equation is ax-x=c
and i saw the problem was solved and the number 1 was in the solution but i don't understand where the 1 came from. can you please explain in really simple steps ? What i got was x=c+x/a

Answer by richard1234(7193)

You want to solve for x in terms of other variables (excluding x). So while your answer x = (c+x)/a is algebraically correct, it isn't completely in terms of other variables.
Factor the LHS to x(a-1) = c, divide both sides by a-1 to get x = c/(a-1).